What is the pOH of a KOH solution that has [OH− ] = 5.81 × 10−3 M?
1
Expert's answer
2020-04-08T13:27:24-0400

Solution:

KOH is a strong base in aqueous media therefore it ionizes 100%.

KOH = K+ + OH


This means that

[OH] = [KOH] = 5.81*10−3 M.

pOH = −log[OH].

Therefore,

pOH = −log[OH] = −log[5.81*10−3] = 2.2358 = 2.24

pOH = 2.24


Answer: 2.24 is the pOH of a KOH solution that has [OH] = 5.81*10−3 M.